EVE AUG-5-1999 shift summary

Alexander Ledovskoy (ledovsko@fnpxsr)
Fri, 6 Aug 1999 00:18:44 -0500

Shift Summary for EVE shift, 5-AUG-99

On Shift: B. Cox, S. Ledovskoy

Hours of beam available at all intensities: 8.0
Hours of beam available at nominal intensity: 8.0
Hours of E832 data (special runs not included): 3.5
Hours of special runs (laser scans, muon runs...):

Accounting of lost time:

HCC Problem/D1 lead failure in accelerator 0.5
Crate 53 problems 3.25
Power glitch 0.25
DDD4 hangs 1.0

========================================================================
Runs ended this shift:
14118 e832_beam
14122 e832_beam

========================================================================
Tapes brought to FCC this shift:
========================================================================
Accesses during shift:
========================================================================
Overview of shift:
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14118 164 15:30 New shift crew takes over smooth data taking.

167 ktevana errors, KFREND fixes it

176 ktevana errors, 3 KFRENDs make triggers running but
pipeline errors prevail. Reset pipeline, kfrend,
everything is OK. 5 spills are lost.

16:00 Vacuum pump oil levels seems to be low (from TV screen)
D.J. is contacted.

16:30 DC online histograms: plane 9, number of hits -
not happy.
CsI RF-L1 timing is on the edge of the window (pipe#)

247 HCC is inefficient error: missing 1563,1561,1453
Reloading HCC does not help.

249 17:30 Beam goes away: D1 lead failure

Jason H., Julie and Jianbo traced HCC problem to bad
CAB_B board. The board was replaced with new one (#9)
HCC probleem is solved.

18:00 Beam is back

18:10 Can not start new run because of crate 53 problems
All usual tricks did not work. See logbook for details.

14122 21:15 Crate 53 problems went away. BEAM_832 started
3h 15m of beam is lost.

21:43 Power glitch. Beam is gone. NM4AN, NM2S2, NM2S3,
NM2SR, NM3S are tripped off. MF13 is OFF also

21:55 All magnets are ON except NM4AN, beam is back
32 This spill is with NM4AN off.
NM4AN is ON.
22:03 MF13 is ON. Everything is in order.

23:00 Triggers stopped. DDD4 veto.
Stop run. DA hangs. Cannot append to tapes
Tapes are out (75% full)
New tapes are in
Preparing for muon run

23:40 muon run INIT hangs.
Elliot traces the problem to plane4 DDD4 stuck
manually reset does not help

========================================================================
Follow up Notes on Previous Shift:
========================================================================
Items to Follow up on Next Shift: